OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -4x - y less than or greater to -2 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

less than or greater to isn't an actual inequality sign is it less than or equal to OR is it greater than or equal to

jimthompson5910 (jim_thompson5910):

this sign \(\large \le\) OR this sign \(\large \ge\)

OpenStudy (anonymous):

First one

jimthompson5910 (jim_thompson5910):

so it is \[\Large -4x - y \le -2\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

we have a solid line since there's a line under the inequality sign

jimthompson5910 (jim_thompson5910):

\[\Large -4x - y \le -2\] \[\Large -4(0) - 0 \le -2\] \[\Large 0 - 0 \le -2\] \[\Large 0 \le -2\] which is FALSE, so (0,0) is NOT in the shaded region

OpenStudy (anonymous):

So its B?

jimthompson5910 (jim_thompson5910):

good